PHOTODISSOCIATION SPECTROSCOPY AND DYNAMICS OF FREE RADICALS \underline{DANIEL M. NEUMARK}, Department of Chemistry, University California, Berkeley, CA 94720, and Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA 94720.